Marvel Parts, Inc., manufactures auto accessories. One of the company’s products is a set of seat covers that can be adjusted to fit nearly any small car. The company has a standard cost system in use for all of its products. According to the standards that have been set for the seat covers, the factory should work 1,030 hours each month to produce 2,060 sets of covers. The standard costs associated with this level of production are: Total Per Set of Covers Direct materials $ 39,140 $ 19.00 Direct labor $ 9,270 4.50 Variable manufacturing overhead (based on direct labor-hours) $ 3,502 1.70 $ 25.20 During August, the factory worked only 640 direct labor-hours and produced 1,600 sets of covers. The following actual costs were recorded during the month: Total Per Set of Covers Direct materials (5,500 yards) $ 29,920 $ 18.70 Direct labor $ 7,520 4.70 Variable manufacturing overhead $ 4,000 2.50 $ 25.90 At standard, each set of covers should require 2.5 yards of material. All of the materials purchased during the month were used in production. Required: 1. Compute the materials price and quantity variances for August. 2. Compute the labor rate and efficiency variances for August. 3. Compute the variable overhead rate and efficiency variances for August. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)
Answer #1
Standard materials price per yard 7.6 =19/2.5
Standard labor hour per unit 0.5 =1030/2060
Standard labor rate per labor hour 9 =4.5/0.5
Standard variable rate per labor hour 3.4 =1.7/0.5
1
Materials price variance 11880 F =29920-(5500*7.6)
Materials quantity variance 11400 U =7.6*(5500-1600*2.5)
2
Labor rate variance 1760 U =7520-(640*9)
Labor efficiency variance 1440 F =9*(640-1600*0.5)
3
Variable overhead rate variance 1824 U =4000-(640*3.4)
Variable overhead efficiency variance 544 F =3.4*(640-1600*0.5)